The following data give the weekly wages of the workers in a firm their total working hours and the mean working hour per workers.

Calculate the mean weekly wage per works.

Wage group (br)

total hour worked

Mean number of hours worked per workers 


800-1000

168

12


1000-1200

170

10


1200-1400

225

3


1400-1600

272

8.5


1600-1800

126

7


1800-2000

91

6.5



1
Expert's answer
2021-10-26T11:37:10-0400

Solution:

Given:

Wage group (br) Class mark (br) total hour worked Mean number of hours worked per workers 


800-1000 900 168 12


1000-1200 1100 170 10


1200-1400 1300 225 3


1400-1600 1500 272 8.5


1600-1800 1700 126 7


1800-2000 1900 91 6.5

Mean weekly wage per worker"=\dfrac{900\times \dfrac{168}{12}+1100\times \dfrac{170}{10}+1300\times \dfrac{225}{3}+1500\times \dfrac{272}{8.5}+1700\times \dfrac{126}{7}+1900\times \dfrac{91}{6.5}}{12+10+3+8.5+7+6.5}"

"=\dfrac{234000}{47}
\\=\$4978.72"